# Solving Nonlinear PDEs with Sparse Radial Basis Function Networks

### v1

@@ 1.2.4. PINNs and other machine learning method for PDEs
Our method is closely related to PINNs [47, 54, 38], which also solve PDEs by minimizing a residual loss over a neural network ansatz. While these works provide convergence results for linear PDEs, they typically rely on large, fixed neural networks without incorporating sparsity or adaptivity, which are central to our approach. Adaptive approaches are developed in [14, 35] for linear PDEs using residual-based neuron enhancement, while our work differs by focusing on nonlinear PDEs and promoting sparsity through RKBS-based regularization. In particular, our framework uses a shallow network trained with a sparsity-promoting $\ell_{1}$ regularized residual loss. This results in a compact representation and better interpretability. Another key difference lies in the optimization strategy. Instead of updating all parameters using gradient-based method, we employ a greedy, boosting-like approach that adaptively selects neurons while maintaining a simple network structure. To further enhance convergence and promote sparsity, we use a second-order optimization scheme. This helps avoid redundant neurons, which are often seen in first-order methods. Although our theory is developed for Gaussian RBF, the approach still applies when we replace RBF with other activation functions commonly used in PINNs.
1.3. Notation, assumptions and examples
@@
(1.10) $\mathcal{R}[u]=0\quad\text{in }\mathcal{F}_{o}.$
@@
Assumption 1.2 (Well-posedness of the PDE).
@@
Such well-posedness properties are well established in the classical theory for several important classes of PDEs; see, for example, [10, 13, 18, 22, 26]. In general, we expect that $\mathcal{V}\subsetneq\mathcal{U}_{o}$. However, a key advantage is that $\mathcal{V}$ is dense in $\mathcal{U}_{o}$. This property follows from the fact that the Gaussian kernel is a “universal kernel”, meaning that finite linear combinations of its shifts can approximate any continuous function on a compact set arbitrarily well [39]. 1.2 together with the density property allows our method to approximate solutions to (1.1). The corresponding analysis is presented in Section 2.3.
@@
Assumption 1.3.
We assume that 1.1 holds. In addition, the functions $\hat{E}(x,l_{E})$ and $\hat{B}(x,l_{B})$ are differentiable in their second arguments, with gradients uniformly bounded in $x\in\overline{D}$ and $x\in\partial D$, respectively.
@@ 1.4. Outline of the paper
2. Theoretical framework
2.1. Integral neural networks and the associated function space
@@
(2.1) $u_{\mu}(x)=\mathcal{N}\mu(x):=\int_{\Omega}\varphi(x;\omega)\mathop{}\!\mathrm{d}\mu(\omega),$
@@
(2.2) $\varphi(x;\omega)=\frac{\sigma^{s}}{\left(\sqrt{2\pi}\sigma\right)^{d}}\exp\left(-\frac{\lVert x-y\rVert^{2}_{2}}{2\sigma^{2}}\right)\quad\text{where }\omega=(y,\sigma)\in\mathbb{R}^{d+1}.$
@@
(2.3) $\varphi(x;\omega)=\sigma^{s-d}\hat{\varphi}\left(\frac{x-y}{\sigma}\right).$
@@
We now define the function space associated with the integral neural networks. Specifically, we consider the set of functions on $D$ that can be written as $\mathcal{N}\mu$ for some signed Radon measure $\mu$. This leads to a reproducing kernel Banach space (RKBS) [4, 34, 66], defined by
(2.4) $\mathcal{V}(D)=\left\{\mathcal{N}\mu:D\to\mathbb{R}\;|\;\mu\in M(\Omega)\right\}$
@@
(2.5) $\lVert u\rVert_{\mathcal{V}(D)}=\inf\{\lVert\mu\rVert_{M(\Omega)}\;|\;\mu\in M(\Omega)\colon\mathcal{N}\mu=u\text{ on }D\}.$
@@
We assume $s\geq d+2+\gamma$ for $\gamma\in(0,1]$ for the rest of this paper. Therefore $\mathcal{V}(D)\hookrightarrow C^{2}(\overline{D})$. By Proposition 2.1, we observe that if the sequence of measures $\mu_{k}$ converges to $\mu$ in the weak-$\ast$ sense, then for any $x\in\overline{D}$,
(2.6) $\partial^{\beta}(\mathcal{N}\mu_{k})(x)=\int_{\Omega}\partial^{\beta}_{x}\varphi(x;\omega)\mathop{}\!\mathrm{d}\mu_{k}(\omega)\to\int_{\Omega}\partial^{\beta}_{x}\varphi(x;\omega)\mathop{}\!\mathrm{d}\mu(\omega)=\partial^{\beta}(\mathcal{N}\mu)(x),$
@@
The result follows directly from the pointwise convergence (2.6) and 1.1. ∎
Theorem 2.5 (Existence of minimizers).
@@
We use direct methods in the calculus of variations to establish the results. Specifically, we prove the existence of minimizers for ($P_{\mu}$), with the result for ($P_{\mu}^{\text{emp}}$) following by a similar argument. Denote $J(\mu)=L(u_{\mu})+\alpha\lVert\mu\rVert_{M(\Omega)}$. Let $\mu^{(k)}\in M(\Omega)$ be a minimizing sequence, i.e., $J(\mu^{(k)})\to\inf_{\mu\in M(\Omega)}J(\mu)$. Then since $\|\mu^{(k)}\|_{M(\Omega)}$ is uniformly bounded, there is a subsequence (without relabelling) that converges to $\bar{\mu}\in M(\Omega)$ in the weak-$\ast$ sense. Now we want to show $J$ is weak-$\ast$ lower semicontinuous such that $\overline{\mu}$ is a minimizer of the problem. First of all, the lower semicontintuity of $\|\cdot\|_{M(\Omega)}$ under weak-$\ast$ convergence is obvious. Second, we observe that for any multi-index $\beta\in\mathbb{N}^{d}$ with $|\beta|\leq 2$, $\partial^{\beta}_{x}\varphi(x;\cdot)\in C_{0}(\Omega)$ for a given $x\in\overline{D}$; see Proposition 2.1. Therefore, by Lemma 2.4 we have $\mathcal{E}[u_{\mu^{(k)}}](x)\to\mathcal{E}[u_{\bar{\mu}}](x)$ $\forall x\in D$, and $\mathcal{B}[u_{\mu^{(k)}}](x)\to\mathcal{B}[u_{\bar{\mu}}](x)$ $\forall x\in\partial D$. Finally, by Fatou’s lemma
@@
2.2. Functional analytical setup and assumptions
Recall the conditions in 1.2 on the well-posedness of the PDEs. We first present an example of a PDE for which these conditions are satisfied. Consider the semilinear Poisson equation in (1.11). For simplicity and without loss of generality, we consider the modified formulation
@@ 2.2. Functional analytical setup and assumptions
A proof of this estimate can be found, for example, in [57, Corollary 1.5]. For the modified equation, under minimal regularity assumptions, we may take $\mathcal{U}=H^{1}_{0}(D)+f_{\mathcal{B}}$ and $\mathcal{F}=H^{-1}(D)\times H^{1/2}(\partial D)$. Indeed, it is not hard to verify that 1.2(2) holds for the pair $(\mathcal{U},\mathcal{F})$ due to the Lipschitz continuity of $\phi_{M}$, and 1.2(3) is satisfied by the monotonicity of the operator $-\Delta u+\phi_{M}(u)$. More precisely, by assuming $f_{\mathcal{E}}\in H^{-1}(D)$ and $f_{\mathcal{B}}\in H^{1/2}(\partial D)$, 1.2(2) follows from the estimates
@@
along with the trace theorems for functions in $H^{1}(D)$. 1.2(3) is justified by
@@
where we have used monotonicity of $\phi_{M}$ and Poincaré inequality for $u-v\in H_{0}^{1}(D)$. Now the choice of $\mathcal{U}_{o}$ and $\mathcal{F}_{o}$ depends on the smoothness of the domain and the data. For example, the classical elliptic regularity for Lipschitz domains allows us to choose $\mathcal{U}_{o}=H^{2}(D)\cap\mathcal{U}$ and $\mathcal{F}_{o}=L^{2}(D)\times H^{3/2}(\partial D)$. Under sufficient additional smoothness assumptions on the domain $D$ and the data $f_{\mathcal{E}}$ and $f_{\mathcal{B}}$, one may also take $\mathcal{U}_{o}=H^{k}(D)\cap\mathcal{U}$ with $\mathcal{F}_{o}=H^{k-2}(D)\times H^{k-1/2}(\partial D)$ for any integer $k\geq 2$. Recall the embedding $B^{s}_{1,1}\hookrightarrow\mathcal{V}$ established in Appendix C. By the embedding properties of Besov spaces (see, e.g., [59, Proposition 4.6]), for any $k>s$,
@@
Moreover, in certain cases, it is also beneficial to modify the form of the loss function depending on the structure of the underlying PDE. For example, in Dirichlet boundary value problems, incorporating derivatives of the boundary term into the loss function can enhance performance. Such extensions are discussed in the subsequent sections. Related ideas on alternative loss functions can also be found in [1].
2.3. Convergence analysis for the neural network solution
@@
For the rest of this subsection, we further strengthen Assumptions 1.1 and 1.3 by assuming that the functions $\hat{E}(x,l_{E})$ and $\hat{B}(x,l_{B})$ are $C^{0,\gamma}$ in their first variable. By 1.2(1), we denote the unique solution to (1.1) by $u\in\mathcal{U}_{o}$. Let $u^{\ast}=\mathcal{N}\mu^{\ast}\in\mathcal{V}$ be an approximation of $u$ in $\mathcal{U}_{o}$. By the continuity of $\mathcal{R}|_{\mathcal{U}_{o}}$ in 1.2(2), it follows that
@@
Theorem 2.9.
Let $u\in\mathcal{U}$ denote the unique solution to (1.1) and $u^{\ast}=\mathcal{N}\mu^{\ast}\in\mathcal{V}$. For a given regularization parameter $\alpha>0$ and collocation number $K=K_{1}+K_{2}>0$, let $u_{K,\alpha}=\mathcal{N}\mu_{K,\alpha}$ be a global minimizer of the empirical problem ($P_{\mu}^{\text{emp}}$). Assume that Assumptions 2.8 and 2.6 hold. As $K_{1},K_{2}\to\infty$, any limit point $\mu_{\alpha}$ of $\mu_{K,\alpha}$ in $M(\Omega)$ defines a function $u_{\alpha}=\mathcal{N}\mu_{\alpha}$ that satisfies
@@
if 2.6(2) holds, where $C>0$ is a generic constant independent of $\alpha$, $\mu^{\ast}$, $u$ and $u^{\ast}$. Moreover, the convergence $u_{K,\alpha}\to u_{\alpha}$ (up to a subsequence) holds in $C^{1}(\overline{\Omega})$.
Remark 2.10.
Theorem 2.9 yields the following implications.
@@
    Using the above estimate, one can select the appropriate $\alpha$ such that $u_{\alpha}$ converges to the true solution $u$ in $\mathcal{U}$. In the first case, by density of $\mathcal{V}$ in $\mathcal{U}_{0}$, one can select $u^{\ast}=u^{\delta}\in\mathcal{V}$ such that $\|u-u^{\delta}\|_{\mathcal{U}_{o}}^{2}=\delta$. Let $\alpha=\min(\delta/\lVert u^{\delta}\rVert_{\mathcal{V}},\delta)$, we observe that $\alpha\to 0$ as $\delta\to 0$. Then, by applying the estimate from Theorem 2.9.
@@
We now prove Theorem 2.9. Denote $\hat{J}(\mu)=\hat{L}(\mathcal{N}\mu)+\alpha\lVert\mu\rVert_{M(\Omega)}$. Since $u_{K,\alpha}=\mathcal{N}\mu_{K,\alpha}$ is a global minimizer of ($P_{\mu}^{\text{emp}}$), it satisfies
@@
Now we show the estimate for $u_{\alpha}$. By Lemma 2.4, we have $\mathcal{E}[u_{K,\alpha}](x)\to\mathcal{E}[u_{\alpha}](x)$ $\forall x\in D$, and $\mathcal{B}[u_{K,\alpha}](x)\to\mathcal{B}[u_{\alpha}](x)$ $\forall x\in\partial D$. Furthermore, applying Fatou’s lemma for varying measures [50, p. 269] and using 2.8, we obtain
@@
(2.7) $\|\mathcal{R}[u_{\alpha}]\|_{L^{2}_{\lambda}}^{2}\leq\|\mathcal{R}[u^{\ast}]\|_{L^{2}_{\lambda}}^{2}+\alpha\|\mu^{\ast}\|_{M(\Omega)}.$
We now utilize Equation 2.7 to show the final result. First, we assume 2.6(1) holds. Then, combining this with with 1.2(3), we have
@@
On the other hand, we use 2.6(1) and 1.2(2) to obtain
@@
Together, these inequalities yield the desired result. Now we assume 2.6(2) holds, then the right-hand side of (2.7) is estimated in the same way. For the left-hand side, we use the interpolation equality and 1.2(2),
@@
(2.8) $C(K,\|e\|_{C^{0,\gamma}})^{-1}\lVert e\rVert_{2,\lambda,K}^{2}\leq\lVert e\rVert_{L^{2}_{\lambda}}^{2}\leq C(K,\|e\|_{C^{0,\gamma}})\lVert e\rVert_{2,\lambda,K}^{2},$
@@
Theorem 2.12.
Let $u$ and $u_{K,\alpha}$ denote the functions given in Theorem 2.9. For any $\delta>0$, there exist $u^{\delta}\in\mathcal{V}$ and parameters $\alpha=\alpha(\delta)\to 0$ and $K=K(\delta)\to\infty$ as $\delta\to 0$ such that
@@
if 2.6(2) holds, where $C>0$ is a generic constant.
Remark 2.13.
As noted in Remark 2.10, without further assumptions on the function spaces, the second case guarantees convergence if we know that $\mathcal{V}$ contains the exact solution $u\in\mathcal{U}_{o}$, which ensures that the $\|u^{\delta}\|_{\mathcal{V}}$ remains uniformly bounded. Once again, an alternative remedy is to design a loss function that ensures the corresponding data space is continuously embedded in $\mathcal{F}$. Under such construction, convergence can be established in a similar way without additional assumptions.
@@
Under 2.6(1), it follows that
@@
On the other hand, if 2.6(2) holds, then
@@
2.4. Representer theorem
To further simplify the notation, we introduce the combined collocation set for the residual $\{x_{1,k}\}\cup\{x_{2,k}\}\in\overline{D}=D\cup\partial D$, and the combined set of quadrature weights $\{w_{1,k}\}\cup\{\lambda w_{2,k}\}$. Then we introduce a new global index $k=1,\ldots K_{1}+K_{2}=K$, such that $x_{k}=x_{1,k}$ and $w_{k}=w_{1,k}$ for $k\leq K_{1}$, while for $k>K_{1}$, we set $x_{k}=x_{2,k-K_{1}}$ and $w_{k}=\lambda w_{2,k-K_{1}}$. With this, we introduce the combined linear operator evaluations as

### v2

@@ 1.2.4 PINNs and other machine learning method for PDEs
Our method is closely related to PINNs (Raissi et al., 2019; Shin et al., 2023; Luo and Yang, 2024; Bonito et al., 2025), which also solve PDEs by minimizing a residual loss over a neural network ansatz. While these works provide convergence results for linear PDEs, they typically rely on large, fixed neural networks without incorporating sparsity or adaptivity, which are central to our approach. Adaptive approaches are developed by Cai et al. (2022); Liu and Cai (2022) for linear PDEs using residual-based neuron enhancement, while our work differs by focusing on nonlinear PDEs and promoting sparsity through RKBS-based regularization. Radial basis function network based approaches related to PINNs have been developed by Ramabathiran and Ramachandran (2021); Wang et al. (2023); Bai et al. (2023). In particular, similar to our framework, Wang et al. (2023) propose a shallow RBF network trained with a sparsity-promoting $\ell_{1}$ regularized residual loss to obtain compact representation and better interpretability. However, a scaling of the feature functions in (1.3) that allows for stable computation of second derivatives for arbitrarily small length scale $\sigma>0$ has not been employed. This is a prerequisite for the analysis in terms of RKBS provided in our work, which enables theoretical guarantees in terms of generalization (error estimates), and adaptive training methods that do not require a large initial choice of the network width.
A key difference of our method compared to other approaches using (regularized) RBF networks is the optimization strategy. Instead of updating all parameters using gradient-based method, we employ a greedy, boosting-like approach that adaptively selects neurons while maintaining a simple network structure. To further enhance convergence and promote sparsity, we use a second-order optimization scheme. This helps avoid redundant neurons, which are often seen in first-order methods. Although our theory is developed for Gaussian RBF, the approach still applies when we replace RBF with other activation functions commonly used in PINNs. Other related approaches, such as variational or weak formulation of loss functions (Yu et al., 2018; Zang et al., 2020), may be explored in the future.
It has been widely observed that the performance of PINNs is highly sensitive to the choice of penalty parameter $\lambda$ for enforcing boundary conditions (Wang et al., 2021, 2022). As our method employs a similar loss function, it inherits this sensitivity as well. This motivates alternative treatment of boundary conditions, which will be discussed later in Section 3.4 and 4.1.3. Related to this, a linear elliptic problem is studied by Bonito et al. (2025) for a class of collocation methods including PINNs formulations. Using the theory of optimal recovery and regularity theory in Besov scales, an improved PINNs loss is provided that allows for better theoretical guarantees when minimizing the loss over a (neural network) model class that approximates functions in Besov spaces. In our work, we directly regularize in a variation norm, which is related to a Besov norm, see Section 2.1, and we make similar observations on the practical and theoretical deficiency of the standard PINNs loss.
1.3 Notation, assumptions and examples
@@
$\mathcal{R}[u]=0\quad\text{in }\mathcal{F}_{o}.$ (1.10)
@@
Assumption 2 (Well-posedness of the PDE).
@@
Such well-posedness properties are well established in the classical theory for several important classes of PDEs; see, for example, Brezis (2011); Caffarelli and Cabré (1995); Dafermos (2005); Evans (2022); Gilbarg and Trudinger (1977). In general, we expect that $\mathcal{V}\subsetneq\mathcal{U}_{o}$. However, a key advantage is that $\mathcal{V}$ is dense in $\mathcal{U}_{o}$. This property follows from the fact that the Gaussian kernel is a “universal kernel”, meaning that finite linear combinations of its shifts can approximate any continuous function on a compact set arbitrarily well (Micchelli et al., 2006). 2 together with the density property allows our method to approximate solutions to (1.1). The corresponding analysis is presented in Section 2.3.
@@
Assumption 3.
We assume that 1 holds. In addition, the functions $\hat{E}(x,l_{E})$ and $\hat{B}(x,l_{B})$ are differentiable in their second arguments, with gradients uniformly bounded in $x\in\overline{D}$ and $x\in\partial D$, respectively.
@@ 1.3.1 Notation summary
    • $H^{s}(A)=W^{s,2}(A)$: Sobolev space with smoothness index $s$ and integrability index $2$.
    • $B^{s}_{p,q}(A)$: Besov space with smoothness index $s$, integrability $p$, and summability $q$.
@@ 1.4 Outline of the paper
2 Theoretical framework
2.1 Integral neural networks and the associated function space
@@
$u_{\mu}(x)=\mathcal{N}\mu(x):=\int_{\Omega}\varphi(x;\omega)\mathop{}\!\mathrm{d}\mu(\omega),$ (2.1)
@@
$\varphi(x;\omega)=\frac{\sigma^{s}}{\left(\sqrt{2\pi}\sigma\right)^{d}}\exp\left(-\frac{\lVert x-y\rVert^{2}_{2}}{2\sigma^{2}}\right)\quad\text{where }\omega=(y,\sigma)\in\mathbb{R}^{d+1}.$ (2.2)
@@
$\varphi(x;\omega)=\sigma^{s-d}\hat{\varphi}\left(\frac{x-y}{\sigma}\right).$ (2.3)
@@
We now define the function space associated with the integral neural networks. Specifically, we consider the set of functions on $D$ that can be written as $\mathcal{N}\mu$ for some signed Radon measure $\mu$. This leads to a reproducing kernel Banach space (RKBS) (Bartolucci et al., 2023; Lin et al., 2022; Zhang et al., 2009), defined by
$\mathcal{V}(D)=\left\{\mathcal{N}\mu:D\to\mathbb{R}\;|\;\mu\in M(\Omega)\right\}$ (2.4)
@@
$\lVert u\rVert_{\mathcal{V}(D)}=\inf\{\lVert\mu\rVert_{M(\Omega)}\;|\;\mu\in M(\Omega)\colon\mathcal{N}\mu=u\text{ on }D\}.$ (2.5)
@@
We assume $s\geq d+2+\gamma$ for $\gamma\in(0,1]$ for the rest of this paper. Therefore $\mathcal{V}(D)\hookrightarrow C^{2}(\overline{D})$. By Proposition 4, we observe that if the sequence of measures $\mu_{k}$ converges to $\mu$ in the weak-$\ast$ sense, then for any $x\in\overline{D}$,
$\partial^{\beta}(\mathcal{N}\mu_{k})(x)=\int_{\Omega}\partial^{\beta}_{x}\varphi(x;\omega)\mathop{}\!\mathrm{d}\mu_{k}(\omega)\to\int_{\Omega}\partial^{\beta}_{x}\varphi(x;\omega)\mathop{}\!\mathrm{d}\mu(\omega)=\partial^{\beta}(\mathcal{N}\mu)(x),$ (2.6)
@@
The result follows directly from the pointwise convergence (2.6) and 1. ∎
Theorem 8 (Existence of minimizers).
@@
We use direct methods in the calculus of variations to establish the results. Specifically, we prove the existence of minimizers for ($P_{\mu}$), with the result for ($P_{\mu}^{\text{emp}}$) following by a similar argument. Denote $J(\mu)=L(u_{\mu})+\alpha\lVert\mu\rVert_{M(\Omega)}$. Let $\mu^{(k)}\in M(\Omega)$ be a minimizing sequence, i.e., $J(\mu^{(k)})\to\inf_{\mu\in M(\Omega)}J(\mu)$. Then since $\|\mu^{(k)}\|_{M(\Omega)}$ is uniformly bounded, there is a subsequence (without relabelling) that converges to $\bar{\mu}\in M(\Omega)$ in the weak-$\ast$ sense. Now we want to show $J$ is weak-$\ast$ lower semicontinuous such that $\overline{\mu}$ is a minimizer of the problem. First of all, the lower semicontintuity of $\|\cdot\|_{M(\Omega)}$ under weak-$\ast$ convergence is obvious. Second, we observe that for any multi-index $\beta\in\mathbb{N}^{d}$ with $|\beta|\leq 2$, $\partial^{\beta}_{x}\varphi(x;\cdot)\in C_{0}(\Omega)$ for a given $x\in\overline{D}$; see Proposition 4. Therefore, by Lemma 7 we have $\mathcal{E}[u_{\mu^{(k)}}](x)\to\mathcal{E}[u_{\bar{\mu}}](x)$ $\forall x\in D$, and $\mathcal{B}[u_{\mu^{(k)}}](x)\to\mathcal{B}[u_{\bar{\mu}}](x)$ $\forall x\in\partial D$. Finally, by Fatou’s lemma
@@
2.2 Functional analytical setup and assumptions
Recall the conditions in 2 on the well-posedness of the PDEs. We first present an example of a PDE for which these conditions are satisfied. Consider the semilinear Poisson equation in (1.11). For simplicity and without loss of generality, we consider the modified formulation
@@ 2.2 Functional analytical setup and assumptions
A proof of this estimate can be found, for example, in (Taylor, 1996, Corollary 1.5). For the modified equation, under minimal regularity assumptions, we may take $\mathcal{U}=H^{1}_{0}(D)+f_{\mathcal{B}}$ and $\mathcal{F}=H^{-1}(D)\times H^{1/2}(\partial D)$. Indeed, it is not hard to verify that 2(2) holds for the pair $(\mathcal{U},\mathcal{F})$ due to the Lipschitz continuity of $\phi_{M}$, and 2(3) is satisfied by the monotonicity of the operator $-\Delta u+\phi_{M}(u)$. More precisely, by assuming $f_{\mathcal{E}}\in H^{-1}(D)$ and $f_{\mathcal{B}}\in H^{1/2}(\partial D)$, 2(2) follows from the estimates
@@
along with the trace theorems for functions in $H^{1}(D)$. 2(3) is justified by
@@
where we have used monotonicity of $\phi_{M}$ and Poincaré inequality for $u-v\in H_{0}^{1}(D)$. Now the choice of $\mathcal{U}_{o}$ and $\mathcal{F}_{o}$ depends on the smoothness of the domain and the data. For example, the classical elliptic regularity for Lipschitz domains allows us to choose $\mathcal{U}_{o}=H^{2}(D)\cap\mathcal{U}$ and $\mathcal{F}_{o}=L^{2}(D)\times H^{3/2}(\partial D)$. Under sufficient additional smoothness assumptions on the domain $D$ and the data $f_{\mathcal{E}}$ and $f_{\mathcal{B}}$, one may also take $\mathcal{U}_{o}=H^{k}(D)\cap\mathcal{U}$ with $\mathcal{F}_{o}=H^{k-2}(D)\times H^{k-1/2}(\partial D)$ for any integer $k\geq 2$. Recall the embedding $B^{s}_{1,1}\hookrightarrow\mathcal{V}$ established in Appendix C. By the embedding properties of Besov spaces (see, e.g., Triebel, 2006, Proposition 4.6), for any $k>s$,
@@
Moreover, in certain cases, it is also beneficial to modify the form of the loss function depending on the structure of the underlying PDE. For example, in Dirichlet boundary value problems, incorporating derivatives of the boundary term into the loss function can enhance performance. Such extensions are discussed in the subsequent sections. Related ideas on alternative loss functions can also be found in Ainsworth and Dong (2024).
2.3 Convergence analysis for the neural network solution
@@
For the rest of this subsection, we further strengthen Assumptions 1 and 3 by assuming that the functions $\hat{E}(x,l_{E})$ and $\hat{B}(x,l_{B})$ are $C^{0,\gamma}$ in their first variable. By 2(1), we denote the unique solution to (1.1) by $u\in\mathcal{U}_{o}$. Let $u^{\ast}=\mathcal{N}\mu^{\ast}\in\mathcal{V}$ be an approximation of $u$ in $\mathcal{U}_{o}$. By the continuity of $\mathcal{R}|_{\mathcal{U}_{o}}$ in 2(2), it follows that
@@
Theorem 12.
Let $u\in\mathcal{U}_{o}$ denote the unique solution to (1.1) and $u^{\ast}=\mathcal{N}\mu^{\ast}\in\mathcal{V}$. For a given regularization parameter $\alpha>0$ and collocation number $K=K_{1}+K_{2}>0$, let $u_{K,\alpha}=\mathcal{N}\mu_{K,\alpha}$ be a global minimizer of the empirical problem ($P_{\mu}^{\text{emp}}$). Assume that Assumptions 11 and 9 hold. As $K_{1},K_{2}\to\infty$, any limit point $\mu_{\alpha}$ of $\mu_{K,\alpha}$ in $M(\Omega)$ defines a function $u_{\alpha}=\mathcal{N}\mu_{\alpha}$ that satisfies
@@
if 9(2) holds, where $C>0$ is a generic constant independent of $\alpha$, $\mu^{\ast}$, $u$ and $u^{\ast}$. Moreover, the convergence $u_{K,\alpha}\to u_{\alpha}$ (up to a subsequence) holds in $C^{1}(\overline{\Omega})$.
Remark 13.
Theorem 12 yields the following implications.
@@
    Using the above estimate, one can select the appropriate $\alpha$ such that $u_{\alpha}$ converges to the true solution $u$ in $\mathcal{U}$. In the first case, by density of $\mathcal{V}$ in $\mathcal{U}_{0}$, one can select $u^{\ast}=u^{\delta}\in\mathcal{V}$ such that $\|u-u^{\delta}\|_{\mathcal{U}_{o}}^{2}=\delta$. Let $\alpha=\min(\delta/\lVert u^{\delta}\rVert_{\mathcal{V}},\delta)$, we observe that $\alpha\to 0$ as $\delta\to 0$. Then, by applying the estimate from Theorem 12.
@@
We now prove Theorem 12. Denote $\hat{J}(\mu)=\hat{L}(\mathcal{N}\mu)+\alpha\lVert\mu\rVert_{M(\Omega)}$. Since $u_{K,\alpha}=\mathcal{N}\mu_{K,\alpha}$ is a global minimizer of ($P_{\mu}^{\text{emp}}$), it satisfies
@@
Now we show the estimate for $u_{\alpha}$. By Lemma 7, we have $\mathcal{E}[u_{K,\alpha}](x)\to\mathcal{E}[u_{\alpha}](x)$ $\forall x\in D$, and $\mathcal{B}[u_{K,\alpha}](x)\to\mathcal{B}[u_{\alpha}](x)$ $\forall x\in\partial D$. Furthermore, applying Fatou’s lemma for varying measures (Royden, 1988, p. 269) and using 11, we obtain
@@
$\|\mathcal{R}[u_{\alpha}]\|_{L^{2}_{\lambda}}^{2}\leq\|\mathcal{R}[u^{\ast}]\|_{L^{2}_{\lambda}}^{2}+\alpha\|\mu^{\ast}\|_{M(\Omega)}.$ (2.7)
We now utilize Equation 2.7 to show the final result. First, we assume 9(1) holds. Then, combining this with with 2(3), we have
@@
On the other hand, we use 9(1) and 2(2) to obtain
@@
Together, these inequalities yield the desired result. Now we assume 9(2) holds, then the right-hand side of (2.7) is estimated in the same way. For the left-hand side, we use the interpolation equality and 2(2),
@@
$C(K,\|e\|_{C^{0,\gamma}})^{-1}\lVert e\rVert_{2,\lambda,K}^{2}\leq\lVert e\rVert_{L^{2}_{\lambda}}^{2}\leq C(K,\|e\|_{C^{0,\gamma}})\lVert e\rVert_{2,\lambda,K}^{2},$ (2.8)
@@
Theorem 15.
Let $u$ and $u_{K,\alpha}$ denote the functions given in Theorem 12. For any $\delta>0$, there exist $u^{\delta}\in\mathcal{V}$ and parameters $\alpha=\alpha(\delta)\to 0$ and $K=K(\delta)\to\infty$ as $\delta\to 0$ such that
@@
if 9(2) holds, where $C>0$ is a generic constant.
Remark 16.
As noted in Remark 13, without further assumptions on the function spaces, the second case guarantees convergence if we know that $\mathcal{V}$ contains the exact solution $u\in\mathcal{U}_{o}$, which ensures that the $\|u^{\delta}\|_{\mathcal{V}}$ remains uniformly bounded. Once again, an alternative remedy is to design a loss function that ensures the corresponding data space is continuously embedded in $\mathcal{F}$. Under such construction, convergence can be established in a similar way without additional assumptions.
@@
Under 9(1), it follows that
@@
On the other hand, if 9(2) holds, then
@@
2.4 Representer theorem
To further simplify the notation, we introduce the combined collocation set for the residual $\{x_{1,k}\}\cup\{x_{2,k}\}\in\overline{D}=D\cup\partial D$, and the combined set of quadrature weights $\{w_{1,k}\}\cup\{\lambda w_{2,k}\}$. Then we introduce a new global index $k=1,\ldots,K_{1}+K_{2}=K$, such that $x_{k}=x_{1,k}$ and $w_{k}=w_{1,k}$ for $k\leq K_{1}$, while for $k>K_{1}$, we set $x_{k}=x_{2,k-K_{1}}$ and $w_{k}=\lambda w_{2,k-K_{1}}$. With this, we introduce the combined linear operator evaluations as